OpenStudy (anonymous):

Trigonometric Functions Solve the following equations for 0<=x<=360. cos(70-x)=0.5 @ganeshie8

OpenStudy (anonymous):

The first value is 10 but i'm not sure how to find the second value

ganeshie8 (ganeshie8):

for what angles the cosine function spits out 0.5 ?

OpenStudy (anonymous):

60

ganeshie8 (ganeshie8):

and ?

OpenStudy (anonymous):

is it 300?

ganeshie8 (ganeshie8):

Yep!

ganeshie8 (ganeshie8):

therefore 70-x can equal 60 or 300

ganeshie8 (ganeshie8):

70-x = 60 solving x gives you x = 10

ganeshie8 (ganeshie8):

70 - x = 300 solving x gives you -230

ganeshie8 (ganeshie8):

Recall that the period of cosine is 360, so adding 360 to the angle wont change the value of cosine

ganeshie8 (ganeshie8):

so add 360 to -230 to make it positive

ganeshie8 (ganeshie8):

-230 + 360 = ?

OpenStudy (anonymous):

130

OpenStudy (anonymous):

Thank you @ganeshie8 :)

ganeshie8 (ganeshie8):

yw